Home‎ > ‎Reports‎ > ‎Crew100A‎ > ‎

day05


Crew 100A Log Book for February 16, 2011

Photos of the day





Commander's Check-In Report

Ivo Ferreira Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA #4: 17:28 / 18:00

Brief Narrative of Field Mission Results: We continued with the habitability experiments during the morning. Third day with the rover tests in teleoperation. Two of the crew members conducted the third of a series of 12 EVAs to

study the stress related to communication during an EVA. Their heart rate was being monitored to measure the effect of communication in their stress/physical effort. The mission was a total success.

The Parrot UAV flew away with a gust in late afternoon. Search procedure at 19:20 due to weather conditions and low visibility and will resume tomorrow.

EVA Data/Interpretations: N/A (To be completed in the following days)

Engineering/Hab Maintenance: N/A

Plans for Tomorrow: Find the UAV. Install wifi range extender unit and test with rover. Continue with experiments: habitability, rover and drone and EVA to test stress level in the afternoon.

Reports scheduled for today: Engineering, Commander, Journalist and EVA#4

Inventory: None

Support Requested: N/A

Miscellaneous: We just received more general food supplies and gluten-free for crew 101, along with the range extender equipment (high gain antennas, amplifier and cables)



Commander's Report

Ivo Ferreira Reporting

The parrot left the cage.

After a "quick evening nap" we started the day with the already usual procedures for the CASPER experiment (sleep disruption experiment).

Yesterday it was my turn and Ayako's to try the equipment and I am pretty sure there will be some nice sleep deprived data in there.

We spent the morning listening to different sounds and checking our heart rate, pressure and saliva for the habitability experiment while the others set up the EVA experiment and the rover teleoperation for the afternoon.

After lunch it was time to depart on an excursion mission to place some targets for the future EVA missions with ATV, we were happy to finally enjoy some of the beauties of this place.

Today it was time for Kévin and Ayako to do another test for the EVA communication-induced stressed experiment which they completed without any problems, leaving us here at the Hab quite happy.

It was then time for another rover test and some other small trials with the parrot on auto pilot mode. But, all of the sudden, the gusts started and the parrot used the opportunity to leave the cage and run away from us and the rover, I do not know what is the problem with those two but we will try to find it out the next days. It was late in the afternoon, the gusts continued while the sun light disappeared and we decided to stop the rescue mission.

We will restart tomorrow first thing in the morning, let's hope the parrot does not fancy "hide and seek".


Engineering Report

Kévin Theatre Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): 0.82

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.59

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 12 Gallons

Notes/Comments:


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): 12

· Trailer Potable Water Tank Level (inches from bottom): 1/3 Full

· Hab Potable Water Tank Level (inches from bottom): 9

· Potable Water Meter Reading: 421568


Notes/Comments:


GreenHab:

· Condition of plants in Tank 1: green water

· Condition of plants in Tank 2: green things

· Condition of plants in Tank 3: green things

· Greenhab notes/ comments:


Observatory:

· Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

· Used (yes/no): No

· Oil Checked (yes/no): No

· Tire Status: Not checked

· Odometer at end of day: 123611.6

· Notes/Comments on Hab Car:


ATV 1 (Opportunity):

· Used: Yes

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 2 (Spirit):

· Used: No

· Oil Checked: No

· Fuel Consumed:

· Tire Status:

· Comments: Still not fixed


ATV 3 (Viking I):

· Used: Yes

· Oil Checked: No

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 4 (Viking II):

· Used: No

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 70

· Actual temperature upstairs (Farenheit): 68

· Thermostat setting downstairs (Farenheit): 55

· Actual temperature downstairs (Farenheit): 64


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: All is well


General Engineering

1. Notes/Comments:


EVA 4 Report

Kévin Theatre Reporting

EVA #4, 17:00- 18:00

Crew Members: Ivo Ferreira, Ayako Ono, Kévin Théatre

Site Location: 518331 4250490, 0518130 42504448

Transit Mode: foot

Objectives: Mission 2 (testing the influence of stress related to communication on an EVA).

Results: Managed to comply with the 3 critical objectives.

Lesson Learned: Anticipation of problem is very useful in this type of mission.


Journalist Report

Romain Benchenafi Reporting
The wind has been shaking the habitat since this morning and the temperature inside is dropping. Nevertheless the morale is still up, we even managed to find the time having fun between two experiments. Unfortunately a gust just took one of our drones away, we hope we will find it tomorrow.